OpenStudy (anonymous):

express as a single logarithm and, if possible,simplify log(a) 32+log(a) 4

OpenStudy (anonymous):

log(a) 32+log(a) 4 = log(a) (32x4) = log(a) 128

OpenStudy (anonymous):

Apply the log rule log(xy) = logx + logy

OpenStudy (anonymous):

Thank You i got

OpenStudy (anonymous):

it

OpenStudy (kainui):

Remember, logs and exponents have the same rules.

OpenStudy (anonymous):

@Stunna1635 you are welcome :) @Kainui Trying to get rid of the poop colour, eh ? ;)

OpenStudy (anonymous):

Given that log (a)7=1.946, and log(a)5= 1.609, find log(a) (7 over 5) 7/5

OpenStudy (anonymous):

Stunna you have to rmb your log rules ... for this its basically the Quotient Rule of Logarithms .... log (a) (x/y) = log (a) x -log (a) y

OpenStudy (anonymous):

log (a) (7/5) = log (a) 7 - log (a) 5 Put in the values given and you will get your answer.

OpenStudy (anonymous):

Stunna do you not have a formula list to refer to ?

OpenStudy (anonymous):

http://www.eeweb.com/toolbox/math-help Here you go ....

OpenStudy (anonymous):

okay, and no i will google it i did not get textbook

OpenStudy (anonymous):

Thank You again

OpenStudy (anonymous):

pauls online math notes has cheat sheets for algebra, trig, calculus, and i think differential equations

OpenStudy (anonymous):